### v5

Abstract
This paper investigates preconditioned conjugate gradient techniques for solving kernel ridge regression (KRR) problems with a medium to large number of data points ($10^{4}\leq N\leq 10^{7}$), and it describes two methods with the strongest guarantees available. The first method, RPCholesky preconditioning, accurately solves the full-data KRR problem in $\mathcal{O}(N^{2})$ arithmetic operations, assuming sufficiently rapid polynomial decay of the kernel matrix eigenvalues. The second method, KRILL preconditioning, offers an accurate solution to a restricted version of the KRR problem involving $k\ll N$ selected data centers at a cost of $\mathcal{O}((N+k^{2})k\log k)$ operations. The proposed methods efficiently solve a range of KRR problems, making them well-suited for practical applications.
1 Motivation
Kernel ridge regression (KRR) is a machine learning method for approximating an unknown function $y=f(\bm{x})$ from a set of input–output pairs
$(\bm{x}^{(1)},y_{1}),(\bm{x}^{(2)},y_{2}),\ldots,(\bm{x}^{(N)},y_{N})\in\mathcal{X}\times\mathbb{R},$
where the inputs take values in an arbitrary set $\mathcal{X}$. KRR is based on a positive-definite kernel $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ that captures the “similarity” between a pair of inputs. KRR produces a nonlinear prediction function $\widehat{f}:\mathcal{X}\to\mathbb{R}$ of the form
$\widehat{f}(\bm{x})=\sum_{i=1}^{N}\beta_{i}K(\bm{x}^{(i)},\bm{x})$
that models the input–output relationship. KRR is widely used in data science and scientific computing, for example, when predicting the chemical properties of molecules [20, 58]. Section˜2 provides more details about KRR.
In practice, KRR is often limited to small- or medium-sized data sets ($N\leq 10^{4}$) because the computation time can grow rapidly with the number of data points. KRR involves linear algebra computations using the $N\times N$ kernel matrix $\bm{A}$ with entries $a_{ij}=K(\bm{x}^{(i)},\bm{x}^{(j)})$. To obtain the prediction function $\widehat{f}$, we must solve an $N\times N$ positive-definite linear system, which requires $\mathcal{O}(N^{3})$ arithmetic operations using the standard direct method based on Cholesky decomposition. Because of this poor computational scaling, survey articles on machine learning for scientific audiences, such as [55], suggest using KRR for small data sets and neural networks for larger data sets. Reliably scaling KRR to larger data sets with $N\geq 10^{4}$ data points remains an open research challenge.
To improve the scalability of KRR, two approaches based on preconditioned conjugate gradient (CG) have been developed. The first approach aims to solve the full-data KRR problem in $\mathcal{O}(N^{2})$ operations [17, 5]. The second approach restricts the KRR problem to $k\ll N$ centers chosen from the data points [52] and aims to solve the restricted problem in $\mathcal{O}(kN)$ operations [47]. However, due to the ill-conditioned nature of many KRR problems, CG-based approaches can only succeed when there is a high-quality preconditioner. Much of the research on solving KRR problems over the past decade has focused on proposing preconditioners [17, 47, 26, 48, 25, 51], empirically testing preconditioners [17, 51, 8], and theoretically analyzing preconditioners [5, 47, 26, 48, 25]. Despite this extensive body of work, a user of KRR might struggle to identify the best preconditioner, since many ideas have been proposed and there has not been a recent comparison.
The primary purpose of a preconditioner is to minimize the number of CG iterations required to solve a linear system to a desired level of accuracy. Each CG iteration can be computationally expensive as it involves either one multiplication with the $N\times N$ kernel matrix (for full-data KRR) or two multiplications with the $N\times k$ kernel submatrix (for restricted KRR). Additionally, the kernel matrix or submatrix may not fit in working memory, in which case kernel matrix entries must be accessed from storage or regenerated on the fly at each CG iteration, which adds to the computational cost. This paper seeks to reduce the expense of KRR by identifying preconditioners that are both efficient to construct and capable of controlling the number of CG iterations required for convergence.
To be useful in practice, a preconditioner must work reliably for a range of KRR instances, and it must be robust against adverse conditions (such as poorly conditioned kernel matrices). Additionally, the preconditioner must not require delicate parameter tuning. Such a robust and no-hassle preconditioner is vital for scientific software that can handle KRR problems framed by different users. Such a preconditioner is critical for cross-validation, as it enables searching for the best kernel model while having confidence that the preconditioner can handle the resulting linear systems efficiently. Unfortunately, as this paper shows, many existing KRR preconditioners are prone to failure when applied to specific KRR problems, limiting their applicability.
To meet the need for higher-quality preconditioners, this paper harnesses randomized matrix approximations [13, 16] that have not been previously applied in the context of KRR but are valuable for achieving robustness. The approximations lead to two randomized preconditioners that allow us to solve more KRR problems in fewer iterations than the techniques in current use [17, 47, 26, 48, 25]. The RPCholesky preconditioner (Algorithm˜1) can be applied to the full-data KRR problem, while the KRILL preconditioner (Algorithm˜2) can be applied to the restricted KRR problem involving $k\ll N$ data centers. Both preconditioners are cheap to implement, are effective for a range of KRR problems, and are supported by strong theoretical guarantees.
RPCholesky.
The RPCholesky preconditioner is designed to solve linear systems involving the full-data $N\times N$ kernel matrix. The RPCholesky algorithm generates a low-rank approximation of the kernel matrix [13], and the preconditioner employs a shift of this low-rank approximation. Any preconditioner based on low-rank approximation, such as RPCholesky, can only be successful when the number of large eigenvalues in the kernel matrix is within the approximation rank. Yet thankfully, many problems in scientific machine learning do exhibit sufficient eigenvalue decay: see Section˜4.1 for a case study drawn from quantum chemistry. Under favorable eigenvalue decay conditions, RPCholesky preconditioning brings the number of CG iterations down to acceptable levels ($10^{0}$ to $10^{2}$).
KRILL.
The KRILL preconditioner is designed to solve restricted KRR problems (Section˜2.2), which are linear systems involving an $N\times k$ submatrix of the full $N\times N$ kernel matrix. KRILL constructs a preconditioner by performing dimensionality reduction on the $N\times k$ kernel submatrix with a sparse random sign embedding [16]. KRILL always reduces the number of CG iterations to acceptable levels ($10^{0}$ to $10^{2}$) regardless of the kernel matrix eigenvalue decay. Thus, KRILL can serve as a robust and reliable tool for large-scale KRR applications, just as krill are a robust and reliable food source in the marine ecosystem.
Broader context.
For decades, preconditioned CG has been the method of choice for solving large-scale linear systems [49] and KRR problems [17, 47, 26, 48, 25]. However, emerging research suggests that row access methods can outperform preconditioned CG for select problems [44, 31]. The relative merits of preconditioned CG versus row access methods are being debated [33, 19], and so far there has not been a complete and definitive comparison between these linear solvers, in part because it is unclear which preconditioned CG and row access methods are most effective. The present work represents partial progress toward a resolution. It demonstrates that RPCholesky and KRILL have the strongest guarantees of any KRR preconditioners; therefore, these methods can serve as exemplars of the preconditioned CG approach in future comparisons.
1.1 Plan for paper
The rest of this paper is organized as follows. Section˜2 presents the RPCholesky and KRILL preconditioning strategies; Section˜3 compares RPCholesky and KRILL to other preconditioners; Section˜4 applies RPCholesky and KRILL to benchmark problems in computational chemistry and physics; Section˜5 presents proofs of the theoretical results; and Section˜6 offers conclusions.
1.2 Notation
For simplicity, we focus on the real setting, although our work extends to complex-valued kernels without significant modification. The transpose and Moore–Penrose pseudoinverse of $\bm{A}$ are denoted $\bm{A}^{*}$ and $\bm{A}^{\dagger}$. Double bars $\left\|\cdot\right\|$ indicate the Euclidean norm of a vector or the spectral norm of a matrix. The matrix condition number is $\kappa(\bm{A})\coloneqq\lVert\bm{A}\rVert\lVert\bm{A}^{-1}\rVert$. For a positive-definite matrix $\bm{M}$, the $\bm{M}$-weighted inner product norm is denoted $\lVert\bm{z}\rVert_{\bm{M}}\coloneqq(\bm{z}^{*}\bm{M}\bm{z})^{1/2}$. The function $\lambda_{i}(\bm{M})$ outputs the $i$th largest eigenvalue of $\bm{M}.$ The $\ell_{1}$ norm of the vector $\bm{x}$ is denoted by $\|\bm{x}\|_{\ell_{1}}=\sum_{i}|x_{i}|$. The symbol $|\mathsf{S}|$ denotes the cardinality of a set $\mathsf{S}$.
2 Algorithms and best practices
For the KRR user’s convenience, this early section highlights our recommended approaches for solving KRR problems. Specifically, Section˜2.1 introduces the RPCholesky preconditioner for solving the full-data KRR problem, while Section˜2.2 introduces the KRILL preconditioner for solving the restricted KRR problem.
2.1 Full-data kernel ridge regression
Recall that we are given $N$ input–output data pairs $(\bm{x}^{(1)},y_{1}),(\bm{x}^{(2)},y_{2}),\ldots,(\bm{x}^{(N)},y_{N})$ in $\mathcal{X}\times\mathbb{R}$ for training. Assume that $K:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a positive-definite kernel function, and define the $N\times N$ positive-semidefinite kernel matrix $\bm{A}$ with numerical entries $a_{ij}=K(\bm{x}^{(i)},\bm{x}^{(j)})$.
In full-data kernel ridge regression (KRR), we build a prediction function of the form
$\widehat{f}(\bm{x};\bm{\beta})=\sum_{i=1}^{N}\beta_{i}K(\bm{x}^{(i)},\bm{x}).$
The coefficients $\bm{\beta}\in\mathbb{R}^{N}$ for the function are chosen to minimize the regularized least-squares loss
$L(\bm{\beta})=\lVert\bm{y}-\bm{A}\bm{\beta}\rVert^{2}+\mu\bm{\beta}^{\ast}\bm{A}\bm{\beta}.$
Here, $\mu>0$ is a regularization parameter. Theoretical literature [10] suggests setting the regularization to $\mu=\mathcal{O}(N^{p})$ for a value $p\in(0,2/3)$ depending on the kernel matrix eigenvalue decay and the smoothness of $f$. In practice, $\mu$ is typically chosen via cross-validation or a grid search.
Minimizing $L(\bm{\beta})$ is a quadratic optimization problem whose solution $\bm{\beta}$ satisfies
$(\bm{A}+\mu\mathbf{I})\bm{\beta}=\bm{y}.$ (2.1)
The linear system can be solved in $\mathcal{O}(N^{3})$ time using a Cholesky decomposition of $\bm{A}+\mu\mathbf{I}$ and two triangular solves. However, when there is a medium or large number of data points ($N\geq 10^{4}$), we instead recommend solving (2.1) at a reduced cost using conjugate gradient with RPCholesky preconditioning, described below. We will prove that the strategy based on RPCholesky solves the full-data KRR problem in $\mathcal{O}(N^{2})$ time, provided the kernel matrix eigenvalues decay at a sufficiently fast polynomial rate (Theorem˜2.2).
2.1.1 RPCholesky preconditioning
Algorithm 1 RPCholesky preconditioning
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, right-hand-side vector $\bm{y}\in\real^{N}$, regularization coefficient $\mu$, approximation rank $r$, and tolerance $\varepsilon$.
2:Approximate solution $\bm{\beta}_{\star}$ to $(\bm{A}+\mu\mathbf{I})\bm{\beta}=\bm{y}$.
[…]
5:Define primitives for preconditioned conjugate gradient:
Product $:\bm{\beta}\mapsto\bm{A}\bm{\beta}+\mu\bm{\beta}.$
Preconditioner $:\bm{\beta}\mapsto\bm{U}\left[(\bm{\Sigma}^{2}+\mu\mathbf{I})^{-1}-\mu^{-1}\mathbf{I}\right]\bm{U}^{*}\bm{\beta}+\mu^{-1}\bm{\beta}.$
[…]
To build a preconditioner for the full-data KRR equations (2.1), we begin with a low-rank approximation $\bm{\widehat{A}}$ of the kernel matrix $\bm{A}$. Then, we define the preconditioner
$\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I}.$ (2.2)
There are many ways to obtain a low-rank approximation for this purpose, and we adopt the RPCholesky algorithm (Algorithm˜4, analyzed in [13, 24]).
RPCholesky is a variant of the partial Cholesky decomposition that chooses random pivots at each step according to an evolving probability distribution. For an input parameter $r\in\mathbb{N}$, the algorithm returns a rank-$r$ approximation in factored form:
$\bm{\widehat{A}}=\bm{F}\bm{F}^{\ast}\quad\text{where}\quad\bm{F}\in\real^{N\times r}.$
The matrix $\bm{\widehat{A}}$ is random because it depends on the choice of random pivots. The average error of $\bm{\widehat{A}}$ is comparable to the best rank-$r$ approximation error, as established in [13, Thm. 3.1] and [24, Thm. 4.2]. The algorithm accesses only $(r+1)N$ entries of the kernel matrix; it uses $\mathcal{O}(rN)$ storage; and it expends $\mathcal{O}(r^{2}N)$ arithmetic operations.
We propose an approximation rank $r=\mathcal{O}(\sqrt{N})$, which ensures that we can obtain an eigenvalue decomposition of the preconditioner in $\mathcal{O}(N^{2})$ operations. We then apply preconditioned CG (Algorithm˜3, described in [27, §10.3]) to solve the KRR problem. If the rank $r=\mathcal{O}(\sqrt{N})$ suffices to obtain a good preconditioner, CG terminates in a constant number of iterations and the total operation count for RPCholesky preconditioning is $\mathcal{O}(N^{2})$. The pseudocode for RPCholesky preconditioning is given in Algorithm˜1.
2.1.2 Empirical performance
Here we test the empirical performance of RPCholesky and four other preconditioners across 20 regression and classification problems described in Table˜1 of Appendix˜A. For each problem, we randomly subsample $N=1.5\times 10^{4}$ data points. We standardize the features (subtract the mean, divide by the standard deviation) and measure similarity using the squared exponential kernel
$K(\bm{x},\bm{y})=\exp\bigl(-\tfrac{1}{2\sigma^{2}}\lVert\bm{x}-\bm{y}\rVert^{2}\bigr)\quad\text{with bandwidth $\sigma=\sqrt{d}$.}$ (2.3)
We then formulate each KRR problem (2.1) with a larger regularization parameter $\mu/N=10^{-7}$ (left panel) or a smaller regularization parameter $\mu/N=10^{-8}$ (right panel) and choose the rank $r$ of the preconditioner to be $r=1000$. This choice of parameters is typical for full-data KRR problems [5, 25], and the parameters are applied systematically over the 20 regression and classification problems without additional tuning. Note the smaller value of $\mu$ makes many of these problems highly ill-conditioned. We run 250 CG iterations and declare each KRR instance to be “solved” as soon as the relative residual $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{y}\rVert$ falls below an error tolerance of $\varepsilon=10^{-3}$. This choice of $\varepsilon$ is justified by the fact that the test error plateaus after reaching this tolerance for all the problems we considered.
Larger regularization: $\mu=10^{-7}N$
Smaller regularization: $\mu=10^{-8}N$
Figure 1: Fraction of solved problems versus number of CG iterations for the 20 KRR problem instances in Table˜1.
Figure˜1 charts the fraction of solved problems using RPCholesky and four other preconditioners based on low-rank approximation. Two alternative preconditioners use the partial Cholesky decomposition with greedy pivot selection [26, 57] or uniformly random pivot selection [17, 25]. The final two preconditioners generate randomized low-rank approximations with ridge leverage score sampling (RLS, [1, 39, 48]) or random Fourier features (RFF, [17, 5]). For details of these methods, see Section˜3.1.
RPCholesky is the top-performing method in Figure˜1 with both the larger (left) and the smaller (right) regularization parameters. The closest competitor is RLS, but RPCholesky consistently performs as well or better than RLS. When $\mu/N=10^{-7}$, RPCholesky converges more quickly than RLS on 19 of the 20 test problems and it achieves a speed-up factor of $1.5\times$ for three problems. As another advantage, RPCholesky is simpler to implement and faster to run than RLS, as it eliminates the cost of approximating the ridge leverage scores (see Section˜3.1).
We might worry that RPCholesky changes behavior depending on the random seed, but the algorithm empirically produces results with just $\pm 10\%$ random variation in the number of CG iterations to convergence. To illustrate this point, Figure˜2 shows the distribution of the relative residual $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{y}\rVert$ for the two problems with the fastest (COMET_MC_SAMPLE) and the slowest (w8a) CG convergence. Lines indicate the median error and shaded regions indicate the $20\%$ to $80\%$ error quantiles, which are computed over 100 independent runs. The error quantiles are hardly distinguishable from the median lines, and they are even closer to the median for RPCholesky than the other stochastic methods (uniform, RLS, and RFF).
COMET_MC_SAMPLE Refer to caption
w8a Refer to caption
Figure 2: Relative residual versus number of CG iterations for the problem with the fastest (COMET_MC_SAMPLE, left) and slowest (w8a, right) CG convergence when $\mu/N=10^{-7}$. Note the different vertical axis scales.
The importance of eigenvalue decay is also evident in Figure˜2. In the problem with rapid eigenvalue decay (COMET_MC_SAMPLE, left), RPCholesky converges to the accuracy threshold $\varepsilon=10^{-3}$ in just a single iteration. In the problem with slow eigenvalue decay (w8a, right), RPCholesky converges $120\times$ more slowly. All the methods based on low-rank approximation struggle to control the condition number for the w8a problem. Nonetheless, even for this difficult problem, RPCholesky significantly improves on unpreconditioned CG.
2.1.3 Theoretical guarantees
RPCholesky is guaranteed to solve any full-data KRR problem given sufficiently fast eigenvalue decay in the kernel matrix $\bm{A}$. To explain the eigenvalue decay condition in detail, we introduce a quantitative measure called the $\mu$-tail rank:
Definition 2.1 (Tail rank).
The $\mu$-tail rank of a psd matrix $\bm{A}\in\mathbb{R}^{N\times N}$ is
$\textup{rank}_{\mu}(\bm{A}):=\min\Biggl\{r\geq 0:\,\sum_{i>r}\lambda_{i}(\bm{A})\leq\mu\Biggr\}.$
For example, in KRR problems with the squared exponential kernel (2.3), the $\mu$-tail rank is $\mathcal{O}(\log(1/\mu)^{d})$ where $d$ is the intrinsic dimensionality of the data [2].
Next, we state our main performance guarantee for RPCholesky preconditioning; the proof appears in Section˜5.1.
Theorem 2.2 (RPCholesky preconditioning).
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix, and let $\mu>0$ be a positive number. Construct a random approximation $\bm{\widehat{A}}$ using RPCholesky with block size $B=1$ and approximation rank
$r\geq\textup{rank}_{\mu}(\bm{A})\,\bigl(1+\log\bigl(\operatorname{tr}\bm{A}/\mu\bigr)\bigr).$ (2.4)
With probability at least $1-\delta$, the RPCholesky preconditioner $\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I}$ controls the condition number at a level
$\kappa\bigl(\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}\bigr)\leq 3/\delta.$ (2.5)
Conditional on the event (2.5), when we apply preconditioned CG to the KRR linear system $(\bm{A}+\mu\mathbf{I})\bm{\beta}=\bm{y}$, we obtain an approximation $\bm{\beta}^{(t)}$ to the actual solution $\bm{\beta}$ that satisfies
$\lVert\bm{\beta}^{(t)}-\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}\leq\varepsilon\,\lVert\bm{\beta}\rVert_{\bm{A}+\mu\mathbf{I}}$ (2.6)
[…]
Theorem˜2.2 ensures that RPCholesky-preconditioned CG can solve any full-data KRR problem up to a fixed error tolerance $\varepsilon$ with failure probability $\delta$ when the eigenvalue condition (2.4) is satisfied. This eigenvalue condition ensures that the approximation rank $r$ is large enough to reliably capture the large eigenvalues of the kernel matrix. The factor $\log(\operatorname{tr}\bm{A}/\mu)$ is modest in size, so the expression (2.4) is mainly determined by $\textup{rank}_{\mu}(\bm{A})$, which counts the number of “large” eigenvalues with the rest of the eigenvalues adding up to size $\mu$ or smaller.
Theorem˜2.2 assumes RPCholesky is performed with a block size $B=1$ for simplicity. This theorem can be extended to $B>1$ given a slightly higher approximation rank:
$r\geq\bigl(\textup{rank}_{\mu}(\bm{A})+B\bigr)\,\bigl(1+\log\bigl(\operatorname{tr}\bm{A}/\mu\bigr)\bigr).$
The recent work [24] analyzes RPCholesky with arbitrary block size $B\geq 1$, and it introduces an “accelerated RPCholesky” algorithm that improves the accuracy of block RPCholesky.
If the $\mu$-tail rank is $\mathcal{O}(\sqrt{N})$ and we apply RPCholesky preconditioning with $\mathcal{O}(\sqrt{N})$ columns, Theorem˜2.2 guarantees that we can solve any full-data KRR problem in $\mathcal{O}(N^{2})$ operations. Uniform and greedy Nyström preconditioning admit no similar guarantee (see Section˜3.1). Theorem˜2.2 also provides insight into the case where the $\mu$-tail rank is larger than $\mathcal{O}(\sqrt{N})$. In this case, we may need to use a larger approximation rank $r$, and the construction cost for the preconditioner would exceed $\mathcal{O}(N^{2})$ operations.
Figure˜3 evaluates RPCholesky’s impact on the 20 example problems when $\mu/N=10^{-7}$. Before RPCholesky preconditioning, the condition number $\lambda_{\max}/\lambda_{\min}$ is nearly $10^{7}$ (left panel). RPCholesky reduces the condition number to a level $10^{0}$$10^{4}$ (middle panel). The impact of RPCholesky is similar to the optimal rank-$r$ preconditioner, which eliminates the leading $r$ eigenvalues by setting them equal to $\mu$ (right panel). Yet even the optimal rank-$r$ preconditioner is sensitive to slow kernel matrix eigenvalue decay.
No preconditioner       RPCholesky eigenvalues     Eigenvalues shifted
eigenvalues $\lambda_{i}(\bm{A}+\mu\mathbf{I})$    $\lambda_{i}(\mu\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2})$    by $r$ indices $\lambda_{i+r}(\bm{A}+\mu\mathbf{I})$
Figure 3: Eigenvalue decay for the 20 KRR problems in Table˜1. Left panel shows eigenvalues of the matrix $\bm{A}+\mu\mathbf{I}$ before RPCholesky preconditioning. Middle panel shows eigenvalues of the matrix $\mu\bm{P}^{-1/2}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/2}$ after RPCholesky preconditioning with $r=1000$. Right panel shows eigenvalues $\lambda_{i+r}(\bm{A}+\mu\mathbf{I})$ resulting from the mathematically optimal rank-$r$ preconditioner. Dashed lines indicate the regularization parameter $\mu=10^{-7}N$.
The best application of RPCholesky balances the cost of forming the preconditioner with the cost of the resulting preconditioned CG iterations. As a simple default, we recommend running RPCholesky with $r=10\sqrt{N}$. This default is $1.2\times$ larger than the approximation rank used in Figure˜3 and is large enough to ensure that all 20 test problems are solved in fewer than 120 iterations when $\mu/N=10^{-7}$. See Section˜4.1 for more exploration of the parameter $r$ with a scientific data set.
2.2 Restricted kernel ridge regression
If the number of data points is so large that we cannot apply full-data KRR, we pursue an alternative approach that we call “restricted KRR”, which was proposed in [52]. In restricted KRR, we build a prediction function
$\widehat{f}(\bm{x};\bm{\widehat{\beta}})=\sum_{i=1}^{k}\widehat{\beta}_{i}K(\bm{x}^{(s_{i})},\bm{x}),$
using a subset $\bm{x}^{(s_{1})},\bm{x}^{(s_{2})},\ldots,\bm{x}^{(s_{k})}$ of input points, which are called “centers”. There are many strategies for selecting the centers, such as uniform sampling, ridge leverage score sampling, and RPCholesky. One must balance the computational cost of the center selection procedure against the quality of the centers. For all the experiments in this paper, we use the computationally trivial approach of sampling centers uniformly at random. However, the development of fast procedures for identifying high-quality centers is a topic for future work.
In restricted KRR, the coefficients $\bm{\widehat{\beta}}$ are chosen to minimize the loss function
$L(\bm{\widehat{\beta}})=\lVert\bm{y}-\bm{A}(:,\mathsf{S})\bm{\widehat{\beta}}\rVert^{2}+\mu\bm{\widehat{\beta}}^{\ast}\bm{A}(\mathsf{S},\mathsf{S})\bm{\widehat{\beta}}$,
where $\bm{A}(:,\mathsf{S})$ denotes the submatrix of $\bm{A}$ with column indices in $\mathsf{S}=\{s_{1},\ldots,s_{k}\}$ and $\bm{A}(\mathsf{S},\mathsf{S})$ denotes the submatrix with row and column indices both in $\mathsf{S}$. Minimizing this quadratic loss function leads to a coefficient vector $\bm{\widehat{\beta}}$ that satisfies
$\bigl[\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})\bigr]\,\bm{\widehat{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}.$ (2.7)
Equation (2.7) is a $k\times k$ linear system that is potentially much smaller than the $N\times N$ system (2.1) in full-data KRR. This system can be arbitrarily ill-conditioned and may even be singular. Therefore, to accurately solve restricted KRR problems in finite-precision arithmetic, we add a small multiple of the machine precision to the diagonal entries of the regularizer $\mu\bm{A}(\mathsf{S},\mathsf{S})$.
Solving (2.7) by direct methods is expensive because of the $\mathcal{O}(k^{2}N)$ cost of forming the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. When the number of centers is $k\geq 10^{3}$, we recommend a cheaper approach for solving (2.7) using conjugate gradient with the KRILL preconditioner, described below. We will prove that KRILL preconditioning solves every restricted KRR problem in $\mathcal{O}((N+k^{2})\,k\log k)$ arithmetic operations (Theorem˜2.3).
2.2.1 KRILL preconditioning
Algorithm 2 KRILL preconditioning
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, right-hand-side vector $\bm{y}\in\real^{N}$, regularization coefficient $\mu$, centers $\mathsf{S}$, and tolerance $\varepsilon$.
2:Approximate solution $\widehat{\bm{\beta}}_{\ast}$ to $(\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S}))\bm{\widehat{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}$.
[…]
10:Define primitives for preconditioned conjugate gradient:
Product $:\widehat{\bm{\beta}}\mapsto\bm{A}(:,\mathsf{S})(\bm{A}(\mathsf{S},:)\bm{\widehat{\beta}})+\bm{H}\widehat{\bm{\beta}}$.
[…]
KRILL is based on a randomized approximation of the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$. To form this approximation, we first generate a sparse random sign embedding [33, §9.2]:
$\bm{\Phi}=\tfrac{1}{\sqrt{\zeta}}\begin{bmatrix}\bm{\phi}_{1}&\cdots&\bm{\phi}_{N}\end{bmatrix}\in\real^{d\times N},$
where $\bm{\phi}_{1},\ldots,\bm{\phi}_{N}$ are columns that possess uniform $\pm 1$ values in $\zeta$ uniformly random positions. Next, we construct the matrix $\bm{B}=\bm{\Phi}\bm{A}(:,\mathsf{S})$. Last, to form the KRILL preconditioner, we calculate the approximate Gram matrix $\bm{\widehat{G}}=\bm{B}^{*}\bm{B}$ and add the regularizer $\mu\bm{A}(\mathsf{S},\mathsf{S})$:
$\bm{P}=\bm{B}^{*}\bm{B}+\mu\bm{A}(\mathsf{S},\mathsf{S})=(\bm{\Phi}\bm{A}(:,\mathsf{S}))^{*}(\bm{\Phi}\bm{A}(:,\mathsf{S}))+\mu\bm{A}(\mathsf{S},\mathsf{S}).$ (2.8)
The pseudocode for KRILL preconditioning is provided in Algorithm˜2.
KRILL preconditioning is efficient because evaluating the approximate Gram matrix $\bm{\widehat{G}}$ requires fewer operations than evaluating the exact Gram matrix $\bm{G}$. The precise number of operations depends on the sparsity $\zeta$ and the embedding dimension $d$. In our theoretical analysis, we assume $\zeta$ and $d$ are functions given in [16] that grow according to $\zeta=\mathcal{O}(\log k)$ and $d=\mathcal{O}(k\log k)$. In practice, we set $\zeta=\lceil\log(k+1)\rceil$ and $d=2k$, which we have found to be sufficient to obtain fast CG convergence on all instances. (Smaller values for the sparsity parameter, such as $\zeta=4$, seems to be reliable as well [9].) Recent research [54, 14] has improved the theoretical understanding of sparse random sign embeddings, but the optimal sparsity for a given embedding dimension remains uncertain. After including the cost of the CG iterations, the total operating cost of KRILL preconditioning is therefore $\mathcal{O}((N+k^{2})\,k\log k)$ in theory and $\mathcal{O}(Nk\log k+k^{3})$ in practice.
2.2.2 Empirical performance
Figure˜4 evaluates the performance of KRILL across the 20 regression and classification problems described in Table˜1. For each problem, we subsample $N=4\times 10^{4}$ data points and select $k=1000$ centers uniformly at random. We standardize the features and apply KRR using the squared exponential kernel (2.3) with the bandwidth $\sigma=\sqrt{d}$. We either set the regularization to a large value $\mu/N=10^{-6}$ or a tiny value $\mu/N=10^{-12}$. We run $100$ CG iterations and declare a problem to be “solved” if the relative residual
$\frac{\lVert\bigl[\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})\bigr]\bm{\widehat{\beta}}-\bm{A}(\mathsf{S},:)\bm{y}\rVert}{\lVert\bm{A}(\mathsf{S},:)\bm{y}\rVert}$ (2.9)
falls below a tolerance of $\varepsilon=10^{-4}$. For comparison, Figure˜4 also evaluates the performance of the FALKON preconditioner that approximates the Gram matrix $\bm{G}$ via Monte Carlo sampling. More details about FALKON are in Section˜3.2.
Large regularization: $\mu=10^{-6}N$
Tiny regularization: $\mu=10^{-12}N$
Figure 4: Fraction of solved problems versus number of CG iterations for the 20 kernel problems in Table˜1.
Examining Figure˜4, we find that KRILL solves all classification and regression problems in 30 or fewer iterations, both when the regularization is large (left) and when the regularization is tiny (right). In contrast, FALKON struggles with the tiny regularization (right panel), and it solves just 8 problems after 30 CG iterations.
KRILL is a randomized algorithm, so its behavior depends on the random seed. Yet the results of KRILL are consistent up to $\pm 20\%$ random variation in the number of CG iterations required for convergence. Figure˜5 shows the median and the 20%–80% quantiles of the relative residual (2.9) for the problems with the fastest (COMET_MC_SAMPLE, left) and slowest (creditcard, right) CG convergence. The quantiles are calculated using 100 independent trials including random subsampling of the data set and random preconditioner formation. The quantiles are tightly concentrated around the median.
COMET_MC_SAMPLE
creditcard
Figure 5: Relative residual versus number of CG iterations for the problems with the fastest (COMET_MC_SAMPLE, left) and slowest (creditcard, right) CG convergence when $\mu/N=10^{-6}$. Thick lines show the median and shaded regions show the $20\%$$80\%$ error quantiles over 100 random trials.
2.2.3 Theoretical guarantees
With the proper parameter choices, KRILL is guaranteed to solve any nonsingular restricted KRR problem. Unlike RPCholesky precondioning, KRILL does not require any eigenvalue decay. Here, we state our main theoretical guarantee, which is proved in Section˜5.2.
Theorem 2.3 (KRILL performance guarantee).
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix, let $\mathsf{S}$ be an index set of cardinality $|\mathsf{S}|=k$, and let $\mu>0$ be a positive number. Define $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ and assume $\bm{M}$ is nonsingular. Generate a sparse random sign embedding $\bm{\Phi}\in\real^{d\times N}$ where the column sparsity $\zeta$ and embedding dimension $d$ are functions of $k$ and $\delta$ given in [16] that satisfy
$\zeta=\mathcal{O}(\log(k/\penalty 50\delta))\quad\text{and}\quad d=\mathcal{O}(k\log(k/\penalty 50\delta)).$ (2.10)
With probability at least $1-\delta$, the KRILL preconditioner $\bm{P}=\bm{A}(\mathsf{S},:)\bm{\Phi}^{\ast}\bm{\Phi}\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ controls the condition number at a level
$\kappa\bigl(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}\bigr)\leq 3.$ (2.11)
Conditional on the event (2.11), when we apply preconditioned CG to the linear system $\bm{M}\widehat{\bm{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}$, we obtain an approximation $\widehat{\bm{\beta}}^{(t)}$ to the actual solution $\widehat{\bm{\beta}}$ that satisfies
$\lVert\widehat{\bm{\beta}}^{(t)}-\widehat{\bm{\beta}}\rVert_{\bm{M}}\leq\varepsilon\lVert\widehat{\bm{\beta}}\rVert_{\bm{M}},$ (2.12)
at any iteration $t\geq\log(2/\penalty 50\varepsilon)$.
Theorem˜2.3 implies that KRILL can solve any nonsingular restricted KRR problem in $\mathcal{O}((N+k^{2})\,k\log k)$ operations. In contrast, FALKON and related preconditioners [47, 34, 48] are only guaranteed to solve restricted KRR problems when the number of centers and the regularization are functions of $k$ satisfying $k=\Omega(\sqrt{N})$ and $\mu=\Omega(\sqrt{N})$ [47, 48]. Given these constraints, it is no surprise that KRILL performs more robustly than FALKON in our experiments.
Figure˜6 evaluates KRILL’s impact on the 20 example problems with a tiny regularization $\mu=10^{-12}N$. Before KRILL preconditioning, the condition number $\lambda_{\max}/\lambda_{\min}$ ranges from $10^{8}$$10^{15}$ (left panel). After KRILL preconditioning, the preconditioned eigenvalues (middle panel) are closely modeled by the inverse square singular values of a $2k\times k$ Gaussian matrix (right panel), leading to a condition number $(1+1/\sqrt{2})^{2}/(1-1/\sqrt{2})^{2}\approx 34.0$. This Gaussian comparison was previously presented in [40, Sec. 2]; and it is an example of the Gaussian universality phenomenon in random matrix theory [54, 14]. The comparison suggests that KRILL performs consistently for all kernel matrices $\bm{A}$, with performance only depending on the embedding dimension $d=2k$.
No preconditioner       KRILL eigenvalues     Eigenvalues $\lambda_{i}(\bm{Z}^{*}\bm{Z})^{-1}$
eigenvalues $\lambda_{i}(\bm{M})$       $\lambda_{i}(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2})$      for $\bm{Z}\sim\mathcal{N}(0,1)^{2k\times k}$
Figure 6: Eigenvalue decay for the 20 KRR problems in Table˜1. Left panel shows eigenvalues of the matrix $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ before KRILL preconditioning (normalized by the largest eigenvalue). Middle panel shows normalized eigenvalues of the matrix $\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}$ after KRILL preconditioning. Right panel shows eigenvalues $\lambda_{i}(\bm{Z}^{*}\bm{Z})^{-1}$ for a Gaussian matrix $\bm{Z}\sim\mathcal{N}(0,1)^{2k\times k}$. Note the different vertical axis scales.
3 Background and comparisons with other preconditioners
In this section, we compare our new conjugate gradient (CG) preconditioners with existing preconditioners for solving the full-data and restricted KRR equations. To begin, we observe that the full-data and restricted KRR equations both take the form
$\bm{M}\bm{z}=\bm{b},$ (3.1)
where $\bm{M}\in\real^{d\times d}$ is a strictly positive definite matrix and $\bm{b}\in\real^{d}$ is a vector. In full-data KRR, the dimension $d$ equals the number of data points, $d=N$. In restricted KRR, the dimension equals the number of data centers, $d=k$.
Preconditioned conjugate gradient.
The CG algorithm [49, §6.7] is a popular approach for solving linear systems of the form (3.1). When $\bm{M}$ is well-conditioned ($\kappa(\bm{M})\approx 1$), CG gives a high-accuracy solution in just a small number of iterations, with each iteration requiring a matrix–vector product with $\bm{M}$. More precisely, the following bound [49, eq. 6.128] controls the convergence rate of the CG iterates $\bm{z}^{(0)},\bm{z}^{(1)},\ldots$ to the solution $\bm{z}$ of the system (3.1):
$\left\|\smash{\bm{z}^{(t)}-\bm{z}}\right\|_{\bm{M}}\leq 2\Biggl(\frac{\sqrt{\kappa(\bm{M})}-1}{\sqrt{\kappa(\bm{M})}+1}\Biggr)^{t}\left\|\smash{\bm{z}^{(0)}-\bm{z}}\right\|_{\bm{M}},$ (3.2)
The convergence is exponentially fast in the $\bm{M}$-norm $\lVert\bm{v}\rVert_{\bm{M}}:=(\bm{v}^{\ast}\bm{M}\bm{v})^{1/2}$. However, the exponential rate depends on the condition number $\kappa(\bm{M})$, and it can be slow when $\kappa(\bm{M})$ is large.
The full-data and restricted KRR equations are typically ill-conditioned, so we need to apply a positive-definite preconditioner $\bm{P}\in\real^{d\times d}$ to these problems to improve the convergence. The preconditioned CG algorithm (Algorithm˜3, described in [27, §10.3]), is equivalent to applying standard CG to the preconditioned system
$\bigl(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}\bigr)\bigl(\bm{P}^{1/2}\bm{z}\bigr)=\bm{P}^{-1/2}\bm{b}.$
The convergence rate of preconditioned CG no longer depends on $\kappa(\bm{M})$, instead depending on $\kappa(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2})$. Fast convergence occurs when $\bm{P}\approx\bm{M}$.
Algorithm 3 Preconditioned conjugate gradient
1:Subroutine Product computing the action $\bm{u}\mapsto\bm{M}\bm{u}$, subroutine Preconditioner computing the action $\bm{u}\mapsto\bm{P}^{-1}\bm{u}$, right-hand-side vector $\bm{b}\in\real^{N}$, and tolerance $\varepsilon$.
2:Approximate solution $\bm{\beta}_{\star}$ to $\bm{M}\bm{\beta}=\bm{b}$.
3:Initialize $\bm{\beta}_{\star}\leftarrow\bm{0},\>\bm{r}\leftarrow\bm{b},\>\bm{z}\leftarrow\texttt{Preconditioner}(\bm{r}),\>\bm{p}\leftarrow\bm{z},$ and $\omega\leftarrow\bm{z}^{*}\bm{r}$
4:while $\|\bm{r}\|\geq\varepsilon\cdot\left\|\bm{b}\right\|$ do
5:  $\bm{v}\leftarrow\texttt{Product}(\bm{p})$
6:  $\eta\leftarrow\omega/\bm{v}^{*}\bm{p}$, $\bm{\beta}_{\star}\leftarrow\bm{\beta}_{\star}+\eta\bm{p}$, $\bm{r}\leftarrow\bm{r}-\eta\bm{v}$
7:  $\bm{z}\leftarrow\texttt{Preconditioner}(\bm{r})$
8:  $\omega_{\rm new}\leftarrow\bm{z}^{*}\bm{r},\gamma\leftarrow\omega_{\rm new}/\omega,\omega\leftarrow\omega_{\rm new}$
9:  $\bm{p}\leftarrow\bm{z}+\gamma\bm{p}$
10:end while
In the rest of this section, we review the history of preconditioned CG methods for the full-data and restricted KRR problems. The section does not cover KRR methods based on sketch-and-solve [6, 1] or row access methods like stochastic gradient descent [18, 32], since the paper’s focus is obtaining the best preconditioned CG method.
3.1 Preconditioners for full-data KRR
To efficiently solve the full-data KRR problem $(\bm{A}+\mu\mathbf{I})\bm{x}=\bm{y}$, we can apply a preconditioner of the form
$\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I},$
where $\bm{\widehat{A}}\in\real^{N\times N}$ is an approximation of the full kernel matrix $\bm{A}$. Ideally, the approximation $\bm{\widehat{A}}\approx\bm{A}$ is highly accurate, and we can quickly apply $\bm{P}^{-1}$ to a vector.
In the KRR literature, $\bm{\widehat{A}}$ is typically constructed as a low-rank approximation [17, 5, 26, 57, 25]. A low-rank approximation can only be accurate when the kernel matrix is close to being low-rank, i.e., when the eigenvalues of $\bm{A}$ decay quickly. However, even under favorable eigenvalue decay conditions, it has been a challenge to identify accurate and computationally tractable low-rank approximations.
One low-rank approximation used in KRR problems is the column Nyström approximation [33, §19.2], which is obtained from a partial Cholesky decomposition with pivoting. The column Nyström approximation takes the form
$\bm{\widehat{A}}=\bm{A}(:,\textsf{S})\bm{A}(\textsf{S},\textsf{S})^{\dagger}\bm{A}(\textsf{S},:),$ (3.3)
where $\textsf{S}=\{s_{1},s_{2},\ldots,s_{r}\}$ is the set of pivots chosen in the Cholesky procedure. The column Nyström approximation is convenient because it is formed from the columns of $\bm{A}$ indexed by S and does not require viewing the rest of the matrix. Yet the approximation accuracy depends on the index set $\mathsf{S}$. In the context of full-data KRR, the most popular strategies for picking the set $\mathsf{S}$ of columns are uniform sampling and greedy selection. Both of these approaches exhibit failure modes.
Failure of uniform sampling.
One simple strategy is to select the column pivots $s_{1},\ldots,s_{r}$ uniformly at random [17, 25]. This uniform sampling method can be effective for some problems, but it fails to explore less populated regions of data space. To illustrate this shortcoming, let $\bm{1}_{m}$ denote the $m$-vector with entries equal to $1$, and consider the kernel matrix
$\bm{A}=\begin{pmatrix}\bm{1}_{N-N^{1/3}}^{\vphantom{*}}\bm{1}_{N-N^{1/3}}^{\ast}&~\\ ~&\bm{1}_{N^{1/3}}^{\vphantom{*}}\bm{1}_{N^{1/3}}^{\ast}\end{pmatrix}$.
Constructing a low-rank approximation for $\bm{A}$ should be easy, since the rank is $2$. However, uniform sampling selects many columns from the left block and neglects columns from the right block that are also needed to ensure the approximation quality. Uniform sampling needs an approximation rank much higher than $2$ and at least as large as $r=\Omega(N^{2/3})$ to build an effective preconditioner.
Failure of greedy selection.
A second strategy for choosing columns is the greedy selection method [26, 57], in which we adaptively select each column pivot $s_{i}$ by finding the largest diagonal element of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after i1i-1 steps of the Cholesky procedure. The greedy method has the opposite failure mode from the uniform method: it focuses on outlier data points and fails to explore highly populated regions of data space. To see this limitation, let $\mathbf{I}_{N^{2/3}}$ denote the $N^{2/3}\times N^{2/3}$ identity matrix, and consider the kernel matrix
$\bm{A}=\bm{1}_{N}^{\vphantom{\ast}}\bm{1}_{N}^{\ast}+\begin{pmatrix}\frac{1}{2N}\,\bm{1}_{N-N^{2/3}}^{\vphantom{*}}\bm{1}_{N-N^{2/3}}^{\ast}&~\\ &\frac{1}{N}\,\mathbf{I}_{N^{2/3}}\end{pmatrix}$.
Constructing a preconditioner for $\bm{A}$ should be easy, since the $\mu$-tail rank (Definition˜2.1) is $\textup{rank}_{\mu}(\bm{A})\leq 2$ for any $\mu\geq N^{-1/3}$. Yet the greedy strategy selects columns from the right block first and misses the left columns. To provide an effective preconditioner, greedy sampling needs an approximation rank $r>N^{2/3}$, which is large enough that all the columns in the right block have been selected.
Algorithm 4 Blocked RPCholesky for psd low-rank approximation
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, approximation rank $k$, block size $B$.
2:Factor matrix $\bm{F}\in\real^{N\times k}$, index set $\mathsf{S}$.
3:Initialize $\bm{F}\leftarrow\bm{0}_{N\times k}$, $\mathsf{S}\leftarrow\emptyset$, $\bm{d}\leftarrow\operatorname{diag}(\bm{A})$, $i\leftarrow 0$
4:while $i<k$ do
5:  Sample iid indices $s_{1},\dots,s_{\min\{B,k-i\}}\sim\bm{d}/\textsc{Sum}(\bm{d})$
6:  $\mathsf{S}^{\prime}\leftarrow\textsc{Unique}(s_{1},\dots,s_{\min\{B,k-i\}})$, $\mathsf{S}\leftarrow\mathsf{S}\cup\mathsf{S}^{\prime}$
7:  $\bm{G}\leftarrow\bm{A}(:,\mathsf{S}^{\prime})-\bm{F}(:,1:i)\bm{F}(\mathsf{S}^{\prime},1:i)^{*}$
8:  $\bm{R}\leftarrow\textsc{Cholesky}(\bm{G}(\mathsf{S}^{\prime},:))$ $\triangleright$ $\bm{G}(\mathsf{S}^{\prime},:)=\bm{R}^{*}\bm{R}$
9:  $\bm{F}\left(:,(i+1):(i+|\mathsf{S}^{\prime}|)\right)\leftarrow\bm{G}\bm{R}^{-1}$
10:  $\bm{d}\leftarrow\bm{d}-\textsc{SquaredRowNorms}\left(\bm{F}\left(:,(i+1):(i+|\mathsf{S}^{\prime}|)\right)\right)$
11:  $i\leftarrow i+|\mathsf{S}^{\prime}|$
12:end while
Advantages of RPCholesky.
In this paper, we propose a new KRR preconditioner that uses RPCholesky (see Algorithm˜4 or [13]) to select the columns for the Nyström approximation. In RPCholesky, we adaptively sample the column pivot $s_{i}$ with probability proportional to the diagonal entries of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after i1i-1 steps of the Cholesky procedure. Because of the adaptive sampling distribution, RPCholesky balances exploration of the small diagonal entries and exploitation of the large diagonal entries, avoiding the failure modes of both the uniform and greedy strategies. RPCholesky is guaranteed to produce an effective preconditioner if the approximation rank $r$ is set to a modest multiple of the $\mu$-tail rank (Theorem˜2.2).
Ridge leverage score sampling.
Ridge leverage score sampling [1, 39, 48] is a more elaborate procedure for selecting the pivots in the Nyström approximation. This method first approximates the ridge leverage scores, which are defined as the diagonal entries $\bm{\ell}^{\lambda}=\operatorname{diag}(\bm{A}(\bm{A}+\lambda\mathbf{I})^{-1})$ for a parameter $\lambda>0$. Then, the method randomly generates a Nyström approximation whose pivots are selected with probabilities proportional to the ridge leverage scores. While RLS sampling can lead to an approximation that is nearly as accurate as RPCholesky (see Figure˜1), it can also produce lower-quality approximations given challenging inputs [13]. Additionally, RLS sampling is slower than the blocked implementation of RPCholesky given in Algorithm˜4, because of the requirement to approximate the ridge leverage scores. For example, when we apply RLS sampling with the RecursiveRLS algorithm [39] to compute a rank-$1000$ approximation of a kernel matrix of size $N=4\times 10^{4}$, the runtime is $2.3\times$ slower than RPCholesky’s runtime.
More expensive Gaussian preconditioner.
A different type of preconditioner for full-data KRR is based on the Gaussian Nyström approximation [25]
$\bm{\widehat{A}}=\bm{A}\bm{Z}(\bm{Z}^{*}\bm{A}\bm{Z})^{\dagger}\bm{Z}^{*}\bm{A},$
where $\bm{Z}\in\real^{N\times r}$ is a matrix with independent standard normal entries. Forming the Gaussian Nyström approximation requires $r$ matrix–vector multiplications with the full kernel matrix. To maintain a cost of $\mathcal{O}(N^{2})$ operations, we can only run Gaussian Nyström with a constant approximation rank $r=\mathcal{O}(1)$, whereas we can run RPCholesky preconditioning with a much larger approximation rank $r=\mathcal{O}(\sqrt{N})$. Because of the higher approximation rank, RPCholesky preconditioning leads to a more accurate approximation with stronger guarantees.
Other approaches.
We briefly mention two other types of approximations that can be used to build preconditioners for the full-data KRR problem. First, certain types of kernel matrices can be approximated using random features [17, 5], but the quality of the approximation tends to be poor. Numerical tests indicate that the random features approach does not yield a competitive preconditioner [17, 59]. See Figures˜1 and 2 for further numerical comparisons.
Second, kernel matrices can be approximated using carefully constructed hierarchical approximations [17, 3, 12, 60] or sparse inverse approximations [50, 61, 28]. These approaches extend beyond low-rank structure and provide richer matrix approximation classes; however, the applicability to high-dimensional data sets remains unclear [3]. So far, these approaches have only been theoretically validated for low-dimensional input data.
3.2 Preconditioners for restricted KRR
To efficiently solve the restricted KRR problem $\bigl[\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})\bigr]\,\bm{\widehat{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}$, we can apply a preconditioner of the form
$\bm{P}=\bm{\widehat{G}}+\mu\bm{A}(\mathsf{S},\mathsf{S}).$ (3.4)
where $\bm{\widehat{G}}\in\real^{k\times k}$ approximates the Gram matrix $\bm{G}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$.
FALKON, the current state-of-the-art.
FALKON-type preconditioners [47, 48, 34] are based on a Monte Carlo approximation of $\bm{G}$. The Monte Carlo approach assumes that each data point $\bm{x}^{(i)}$ is selected as a center with nonzero probability $p_{i}=\mathbb{P}\{i\in\mathsf{S}\}$. The Gram matrix is then approximated as
$\bm{\widehat{G}}=\bm{A}(\mathsf{S},\mathsf{S})\bm{D}(\mathsf{S},\mathsf{S})^{-1}\bm{A}(\mathsf{S},\mathsf{S}),$ (3.5)
where $\bm{D}=\operatorname{diag}(p_{1},p_{2},\ldots,p_{N})$ is the diagonal matrix of the selection probabilities. The matrix entries $g_{ij}$ and $\widehat{g}_{ij}$ can be explicitly written as
$g_{ij}=\sum_{\ell=1}^{N}a_{s_{i},\ell}a_{\ell,s_{j}}\quad\text{and}\quad\widehat{g}_{ij}=\sum_{\ell=1}^{N}\frac{\mathbb{1}\{\ell\in\mathsf{S}\}}{p_{\ell}}a_{s_{i},\ell}a_{\ell,s_{j}}.$
These entries are the same in expectation:
$\sum_{\ell=1}^{N}a_{i,\ell}a_{\ell,j}=\mathbb{E}\Biggl[\sum_{\ell=1}^{N}\frac{\mathbb{1}\{\ell\in\mathsf{S}\}}{p_{\ell}}a_{i,\ell}a_{\ell,j}\Biggr]\quad\text{for fixed indices $1\leq i,j\leq N$,}$
which suggests that the quantities $g_{ij}$ and $\widehat{g}_{ij}$ will be close under appropriate conditions. FALKON [47, 34] is a special case of (3.5) in which the centers are sampled uniformly at random, and FALKON-BLESS [48] is a special case of (3.5) in which the centers are chosen by ridge leverage score sampling.
The main limitation of FALKON-type preconditioners is the Monte Carlo approximation error. To account for the Monte Carlo error, the available analyses [47, 48] assume a large number of centers $k=\Omega(\sqrt{N})$ and a large regularization $\mu=\Omega(\sqrt{N})$. When these assumptions are satisfied, FALKON-type preconditioners solve the restricted KRR equations in $\mathcal{O}((N+k^{2})k)$ operations. However, the assumptions on $k$ and $\mu$ may not be satisfied in practice. For example, $\mu$ is often small when chosen by cross-validation, and [47] contains examples of $\mu$ values up to five orders of magnitude smaller than $\mu=\sqrt{N}$.
Algorithm 5 Sparse sign embedding
1:Embedding dimension $d$, input dimension NN, sparsity level $\zeta$.
2:Embedding matrix $\bm{\Phi}\in\real^{d\times N}$ in sparse (e.g., CSR) format.
[…]
4:for $1\leq j\leq N$ do
5:  $\mathtt{rows}((j-1)\zeta+1:j\zeta)\leftarrow$ distinct uniformly random indices from $\{1,\ldots,d\}$
6:  $\mathtt{cols}((j-1)\zeta+1:j\zeta)\leftarrow j$
7:  $\mathtt{vals}((j-1)\zeta+1:j\zeta)\leftarrow\zeta^{-1/2}\cdot\textsc{Unif}\{\pm 1\}^{\zeta}$
8:end for
[…]
KRILL.
In this paper, we recommend the approximation
$\bm{\widehat{G}}=\bm{A}(\mathsf{S},:)\bm{\Phi}^{\ast}\bm{\Phi}\bm{A}(:,\mathsf{S}),$
where $\bm{\Phi}\in\real^{d\times N}$ is a sparse random sign embedding (see Algorithm˜5 or [16]). With this approximation, we can guarantee the effectiveness of the KRILL preconditioner $\bm{P}=\bm{\widehat{G}}+\mu\bm{A}(\mathsf{S},\mathsf{S})$, regardless of the number of centers $k$ and the regularization $\mu$.
KRILL is inspired by the sketch-and-precondition approach for solving overdetermined least-squares problems [33, §10.5]. In sketch-and-precondition, we use a random embedding to approximate the matrix appearing in the normal equations, and the approximation serves as a preconditioner for solving the least-squares problem. Sketch-and-precondition was first proposed in [45] and later refined in [4, 37, 15, 29, 35]. It has been applied with several different embeddings [33, §§8–9], but the empirical comparisons of [21, Fig. 1] suggest the sparse sign embedding is most efficient.
KRILL differs from classical presentations of sketch-and-precondition because the embedding is applied to the Gram matrix term $\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})$ but not the regularization term $\mu\bm{A}(\mathsf{S},\mathsf{S})$. This idea of “partial sketching” appears in [41, 38], and it improves the accuracy of the approximation. Notwithstanding this difference, KRILL is motivated by the same ideas as classical sketch-and-precondition and its analysis follows the same patterns.
4 Case studies
In this section, we apply our preconditioning strategies to two scientific problems. In Section˜4.1, we apply full-data KRR with RPCholesky preconditioning to predict the chemical properties of a wide range of molecules. In Section˜4.2, we apply restricted KRR with KRILL preconditioning to distinguish exotic particle collisions from a background process.
4.1 HOMO energy prediction
A major goal of chemical machine learning is to search over a large parameter space of molecules and identify candidate molecules which may possess useful properties [20, 58]. To support this goal, scientists have assembled the QM9 data set, which describes the properties of $1.3\times 10^{5}$ organic molecules [43, 46]. Although the QM9 data set is generated by relatively expensive (and highly repetitive) density functional theory calculations, we can use the QM9 data to train machine learning models that efficiently make out-of-sample predictions, limiting the need for density functional theory in the future. Ideally, machine learning would identify a small collection of promising molecules, which scientists could then analyze and test.
A recent journal article [53], which was selected as an editor’s pick in the Journal of Chemical Physics, describes the difficulties in applying KRR to the QM9 data set. The authors aim to predict the highest-occupied-molecular-orbital (HOMO) energy. In their simplest approach, they make HOMO predictions based on the Coulomb matrix representation of molecules, which is a set of features based on the distances between the atomic nuclei. After defining the features, they standardize the features and apply an $\ell_{1}$ Laplace kernel
$K(\bm{x},\bm{x}^{\prime})=\exp\bigl(-\tfrac{1}{\sigma}\lVert\bm{x}-\bm{x}^{\prime}\rVert_{\ell_{1}}\bigr),$
with a bandwidth $\sigma$ and a ridge parameter $\mu$ chosen using cross-validation.
Given the moderate size of the QM9 data set ($N=1.3\times 10^{5}$), it takes just a few seconds to load the data and form individual columns of the kernel matrix. However, on a laptop-scale computer (64GB RAM), there is not enough working memory to store or factorize the complete kernel matrix. The authors of [53] address these computational constraints by randomly subsampling $N=32,000$ or fewer data points and applying full-data KRR to the random sample. In their Fig. 8, they show that the predictive accuracy improves by a factor of $2\times$ as they increase the size of the data from $N=1,000$ to $N=32,000$ data points. Even with $N=32,000$ data points, the predictive accuracy does not saturate, and thus computational constraints are limiting the the predictive accuracy of the KRR model.
Building on this work, we can apply KRR to the complete QM9 data set either using (i) restricted KRR or (ii) full-data KRR with RPCholesky preconditioning. In all our experiments, we split the data into $N=10^{5}$ training points and $M=3\times 10^{4}$ test points, and we measure the predictive accuracy using the symmetric mean absolute percentage error (SMAPE):
$\operatorname{SMAPE}(\bm{\widehat{y}},\bm{y})=\frac{1}{M}\sum_{i=1}^{M}\frac{|\widehat{y}_{i}-y_{i}|}{(|\widehat{y}_{i}|+|y_{i}|)/2}.$ (4.1)
The results are presented in Figures˜7, 8, and 9.
Figure 7: SMAPE versus number of centers / number of subsampled data points for the HOMO energy prediction problem $(N=10^{5})$.
Restricted KRR is a relatively cheap approach for predicting HOMO energies with the QM9 data set, since we can store the $k\times k$ preconditioner and $N\times k$ kernel submatrix in working memory with up to $k=50,000$ centers. Figure˜7 shows that restricted KRR is more accurate than randomly subsampling the data and applying full-data KRR to the random sample. Still, the predictive accuracy increases with the number of centers $k$ and does not saturate even when $k=N/2$.
Full-data KRR is the most accurate approach for predicting HOMO energies, but it relies on repeated matrix–vector multiplications with the full kernel matrix. Since the kernel matrix is too large to store in 64GB of working memory, we can only store one block of the matrix at a time and perform the multiplications in a block-wise fashion. Each multiplication requires tens of minutes of computing, due to the cost of evaluating the kernel matrix each time the entries are needed. The training of the full-data KRR model is very slow unless we find a preconditioner that controls the total number of CG iterations.
Figure 8: (left) Relative residual $\left\|(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\right\|/\left\|\bm{y}\right\|$ and (right) SMAPE for different column Nyström preconditioners for the HOMO energy prediction task
Figure˜8 compares several preconditioning strategies for full-data KRR. The worst strategy is unpreconditioned CG, which leads to no perceivable convergence over the first $100$ iterations. Uniform and greedy Nyström preconditioning (with an approximation rank $r=10^{3}$) improve the convergence speed, but they still require $100$ or more iterations for the predictive accuracy to saturate. RPCholesky (also with $r=10^{3}$) is the fastest method, leading to converged accuracy after $60$ iterations.
Figure 9: Number of CG iterations to achieve a relative residual of $\varepsilon<10^{-3}$ with RPCholesky preconditioning for the HOMO energy prediction task.
Figure˜9 shows that we can reduce the total number of CG iterations even further, by a factor of $5\times$, if we increase the RPCholesky approximation rank from $r=10^{3}$ to $r=10^{4}$. However, we also need to consider the cost of preparing the preconditioner, which requires $\mathcal{O}(Nr^{2})$ operations. To achieve an appropriate balance, we recommend tuning the value of $r$ based on computational constraints, and $r$ should typically be small enough that the preconditioner can be stored in working memory. For example, an intermediate value of $10^{3}<r<10^{4}$ would be most appropriate for the HOMO energy prediction problem given our computing architecture.
Figure˜9 also reveals that RPCholesky performs similarly to uniform pivot selection for a small approximation rank ($r<500$) and performs similarly to greedy pivot selection for a large approximation rank ($r>5000$). It remains an open problem to identify the problems and approximation ranks for which uniform or greedy pivot selection are as successful as RPCholesky.
4.2 Exotic particle detection
The Large Hadron Collider (LHC) is a particle accelerator that is used to search for exotic particles not included in the standard model of physics. Each second, the LHC collides roughly one billion proton–proton pairs and produces one petabyte of data. Yet, exotic particles are believed to be produced in fewer than one per billion collisions, so only a small fraction of the data is relevant. The LHC uses machine learning to identify the most important observations for testing exotic particle models [42].
To help train machine learning algorithms, Baldi, Sadowsky, and Whiteson [7] produced the SUSY data set ($N=5\times 10^{6}$), derived from Monte Carlo simulations of the decay processes of standard model bosons and supersymmetric particles not included in the standard model. Following [47, 48, 30], we will apply restricted KRR to the SUSY data to distinguish between the bosonic and supersymmetric decay processes.
For our KRR application, we split the data into $N=4.5\times 10^{6}$ training data points and $M=5\times 10^{5}$ testing data points, and we select $k=10^{4}$ data centers uniformly at random. Then we standardize the data features and apply a squared exponential kernel (2.3) with bandwidth $\sigma=4$. The parameters $k$ and $\sigma$ are the same ones used in previous work [47, 48], but we use a smaller regularization $\mu/N=2\times 10^{-10}$ (compared to $\mu/N=10^{-6}$ in [47, 48, 30]) which improves the test error from $19.6\%$ to $19.5\%$. We apply KRILL, FALKON, and unpreconditioned CG to solve the restricted KRR equations with these parameter choices. We terminate after forty iterations or when the relative residual falls below $\varepsilon=10^{-4}$, leading to the results in Figures˜10 and 11.
Figure 10: (left) Relative residual $\small\left\|\smash{(\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S}))\,\bm{\widehat{\beta}}-\bm{A}(\mathsf{S},:)\bm{y}}\right\|/\left\|\bm{A}(\mathsf{S},:)\bm{y}\right\|$ and (right) test error for FALKON vs. KRILL for the SUSY classification problem
Figure˜10 demonstrates the superior performance of KRILL, which reaches a test error of $19.5\%$ after just four iterations. Meanwhile, unpreconditioned CG fails to achieve such a low test error after forty iterations, and FALKON converges even more slowly than unpreconditioned CG. One reason for the slow convergence of FALKON is the small regularization parameter $\mu/N=2\times 10^{-10}$ [30, Sec. 3]. Indeed, a previous application [48] of FALKON to the SUSY data set with a larger regularization $\mu/N=10^{-6}$ achieved test error convergence more quickly, in just twenty iterations (still five times slower than KRILL). A major advantage of KRILL is the reliable performance for the full range of $\mu$ values, which empowers users to choose the regularization that gives the lowest test error.
Figure 11: Runtime versus number of centers for the SUSY classification problem ($N=5\times 10^{5}$).
Last, Figure˜11 evaluates the total computation time needed to form the preconditioner and solve the restricted KRR equations up to a relative residual of $\varepsilon=10^{-4}$. For this figure, we randomly subsample $N=5\times 10^{5}$ data points so the kernel submatrix fits in 64GB of working memory, and we compare KRILL, FALKON, and the direct method of forming and inverting $\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$. We find that KRILL is the dominant algorithm once the number of centers reaches $k=1250$. KRILL outperforms the direct method because of the superior $\mathcal{O}(Nk\log k+k^{3})$ computational cost, and KRILL outperforms FALKON because it converges in roughly $4\times$ fewer CG iterations. As an advantage, the direct method solves the restricted KRR equations up to machine precision, whereas KRILL and FALKON only solve the equations up to a relative residual of $\varepsilon=10^{-4}$. Nonetheless, the test error plateaus at a threshold of $\varepsilon=10^{-4}$ in all our experiments.
5 Theoretical results
In this section, we prove our main theoretical results, Theorems˜2.2 and 2.3.
5.1 Proof of Theorem˜2.2
Let $\bm{\widehat{A}}$ be a Nystrom approximation of the psd matrix $\bm{A}$, and set $\bm{P}=\bm{\widehat{A}}+\mu\mathbf{I}$. As the first step of our proof, we will show that
[…]
$2\mathbf{I}\preceq\bm{P}^{-1/\penalty 502}(\bm{A}+\mu\mathbf{I})\bm{P}^{-1/\penalty 502}$
and thus
[…]
Next, a short calculation shows that
[…]
Since the spectral norm is submultiplicative and it equals the largest eigenvalue of a psd matrix, we deduce that
[…]
Combining the bound (5.2) for the minimum eigenvalue with the bound (5.3) for the maximum eigenvalue verifies the condition number bound (5.1).
As our next step, we apply the main RPCholesky error bound [13, Thm. 3.1] with rank $r=\text{rank}_{\mu}(\bm{A})$ and approximation accuracy $1+\varepsilon=2$, which guarantees
$\mathbb{E}\operatorname{tr}\bigl(\bm{A}-\bm{\widehat{A}}\bigr)\leq 2\sum_{i>\text{rank}_{\mu}(\bm{A})}\lambda_{i}(\bm{A})$ (5.4)
for any $r$ satisfying
$r\geq\textup{rank}_{\mu}(\bm{A})\,\bigl(1+\log\bigl(\operatorname{tr}\bm{A}/\mu\bigr)\bigr).$
The right-hand side of (5.4) is bounded by $2\mu$ because of Definition˜2.1 of the $\mu$-tail rank of $\bm{A}$. Hence, combining (5.4) with the condition number bound (5.1) guarantees
[…]
By Markov’s inequality, it follows that
[…]
with failure probability at most $\delta$. We apply the CG error bound (3.2) to obtain
[…]
for each $t\geq 0$, with failure probability at most $\delta$, which is equivalent to the stated convergence guarantee for RPCholesky preconditioning.
5.2 Proof of Theorem˜2.3
Cohen [16, Thm. 4.2] showed that a sparse sign embedding with parameters $d=\mathcal{O}(k\log(k/\delta))$ and $\zeta=\mathcal{O}(\log(k/\delta))$ satisfies the following oblivious subspace embedding property: for any $k$-dimensional subspace $\mathsf{X}\subseteq\real^{N}$, there is probability at least $1-\delta$ that
$(1/2)\left\|\bm{v}\right\|^{2}\leq\left\|\bm{\Phi}\bm{v}\right\|^{2}\leq(3/2)\left\|\bm{v}\right\|^{2}\quad\text{for every $\bm{v}\in\mathsf{X}$}.$ (5.5)
We set $\mathsf{X}=\textup{range}(\bm{A}(:,\mathsf{S}))$ and consider the matrices
$\bm{P}=\bm{A}(\mathsf{S},:)\bm{\Phi}^{\ast}\bm{\Phi}\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})\quad\text{and}\quad\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S}).$
By the subspace embedding property (5.5),
$\lambda_{\rm max}\left(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}\right)$ $=\max_{\bm{z}\neq\bm{0}}\frac{\bm{z}^{*}\bm{M}\bm{z}}{\bm{z}^{*}\bm{P}\bm{z}}\leq\max_{\bm{z}\neq\bm{0}}\;\max\Biggl\{1,\frac{\left\|\bm{A}(:,\mathsf{S})\bm{z}\right\|^{2}}{\left\|\bm{\Phi}\bm{A}(:,\mathsf{S})\bm{z}\right\|^{2}}\Biggr\}\leq 2;$
$\lambda_{\rm min}\left(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}\right)$ $=\min_{\bm{z}\neq\bm{0}}\frac{\bm{z}^{*}\bm{M}\bm{z}}{\bm{z}^{*}\bm{P}\bm{z}}\geq\min_{\bm{z}\neq\bm{0}}\;\min\Biggl\{1,\frac{\left\|\bm{A}(:,\mathsf{S})\bm{z}\right\|^{2}}{\left\|\bm{\Phi}\bm{A}(:,\mathsf{S})\bm{z}\right\|^{2}}\Biggr\}\geq\frac{2}{3}.$
Therefore, $\kappa(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2})\leq 3$, and we apply the CG error bound (3.2) to obtain
$\frac{\lVert\bm{\beta}^{(t)}-\bm{\beta}\rVert_{\bm{M}}}{\lVert\bm{\beta}\rVert_{\bm{M}}}\leq 2\Biggl(\frac{\sqrt{3}-1}{\sqrt{3}+1}\Biggr)^{t}\leq 2\mathrm{e}^{-t}$
for each $t\geq 0$, which is equivalent to the stated performance guarantee.
6 Conclusions
This paper has identified two algorithms for solving kernel ridge regression (KRR) problems with a moderate number of data points ($10^{4}\leq N\leq 10^{7}$). The proposed algorithms based on RPCholesky and KRILL preconditioning are relatively fast, robust, and reliable, outperforming previous approaches [17, 25, 26, 57, 47, 48, 34].
If the number of data points $N$ is not prohibitively large (say, $N\leq 10^{5}$$10^{6}$), we recommend solving the full-data KRR equations using conjugate gradient with RPCholesky preconditioning. The RPCholesky approach enables us solve KRR problems in just $\mathcal{O}(N^{2})$ operations when the kernel matrix eigenvalues decay at a sufficiently fast polynomial rate and the chosen regularization is not too small. For example, when RPCholesky preconditioning is applied to the QM9 data set with $N=1.3\times 10^{5}$ data points, $60$ CG iterations are enough to achieve a high predictive accuracy.
Eigenvalue decay is the main requirement limiting the performance of RPCholesky preconditioning, and this limitation becomes apparent when the regularization $\mu$ is small (see the right panel of Figure˜1). Future work is needed to discover effective preconditioners for KRR problems with slow eigenvalue decay. Methods based on hierarchical decompositions [17, 3, 12, 60] and sparse inverse approximations [50, 28, 61] are promising candidates for this purpose, but there is a concern these methods perform poorly in high dimensions [3].
If the number of data points is so large that $\mathcal{O}(N^{2})$ operations is too expensive, we recommend restricting the KRR equations to $k\ll N$ data centers and solving the restricted KRR equations using conjugate gradient with KRILL preconditioning. The KRILL approach converges to the desired accuracy in just $\mathcal{O}((N+k^{2})k\log k)$ operations for any kernel matrix and regularization parameter $\mu$. For example, when KRILL preconditioning is applied to the SUSY data set with $N=5\times 10^{6}$ data points, CG converges in just $4$ iterations, even with a parameter setting $\mu/N=2\times 10^{-10}$ that would be challenging for other algorithms.
With appropriate parameter choices and neglecting rounding errors, KRILL is guaranteed to control the condition number regardless of the regularization and the kernel matrix eigenvalue decay. Nonetheless, there remain opportunities to improve on KRILL. First, the existing analysis of sparse random sign embeddings [16] does not specify appropriate values of $\zeta$ and $d$ to use in practice. More analysis of sparse sign embeddings is needed to close the theory–practice gap; recent papers [14, 54] make some progress in this direction. A second potential concern with KRILL is numerical stability; see the recent papers [36, 23]. As a practical remedy, we have addressed numerical stability by adding a small shift to the regularizer, but this approach has not been analyzed rigorously. Finally, it may be possible to develop more efficient algorithms for restricted KRR problems when the number of data centers is $k\gg N^{1/2}$. The ideal conjugate gradient algorithm would require $\mathcal{O}(kN)$ operations, but KRILL is more expensive due to the cost of inverting the $k\times k$ preconditioner.
Acknowledgments
We would like to acknowledge helpful conversations with Misha Belkin, Tyler Chen, Jackie Lok, Riley Murray, Madeleine Udell, and Jorge Garza-Vargas.
Appendix A Data sets
In our numerical experiments, we use data from LIBSVM [11], OpenML [56], and UCI [22]. We also use the QM9 dataset [43, 46] which can be found online at https://doi.org/10.6084/m9.figshare.c.978904.v5. See the Github repository https://github.com/eepperly/Fast-Efficient-KRR-Preconditioning for a script to download the data sets.
Table 1: Data sets used in our experiments.
Data set Dimension $d$ Sample size NN Source
Testbed
ACSIncome 11 1 331 600 OpenML
Airlines_DepDelay_1M 9 800 000 OpenML
cod_rna 8 59 535 LIBSVM
COMET_MC_SAMPLE 4 71 712 LIBSVM
connect_4 126 54 045 LIBSVM
covtype_binary 54 464 809 LIBSVM
creditcard 29 227 845 OpenML
diamonds 9 43 152 OpenML
HIGGS 28 500 000 LIBSVM
hls4ml_lhc_jets_hlf 16 664 000 OpenML
ijcnn1 22 49 990 LIBSVM
jannis 54 46 064 OpenML
Medical_Appointment 18 48 971 OpenML
MNIST 784 60 000 OpenML
sensit_vehicle 100 78 823 LIBSVM
sensorless 48 58 509 LIBSVM
volkert 180 46 648 OpenML
w8a 300 49 749 LIBSVM
YearPredictionMSD 90 463 715 LIBSVM
yolanda 100 320 000 OpenML
Additional data sets
QM9 435 133 728 [43, 46]
SUSY 18 5 000 000 UCI